OpenStudy (jcwinner2016):

What Is (3/4)-2 Also Use A Formula

OpenStudy (3mar):

Can you subtract rational numbers?

OpenStudy (lannyxx):

use a formula for what?

OpenStudy (jcwinner2016):

To Know How To Use In Other Questions

OpenStudy (lannyxx):

Oh, you want to know the procedure of solving those problems, y'mean?

OpenStudy (jcwinner2016):

Yes

OpenStudy (3mar):

Let's simplify: \[(3/4)-2=\frac{ 3 }{ 4 }-2\] it is important to use parentheses.

OpenStudy (3mar):

then what is the denominator of 2? Is not it \[\frac{ 2 }{ 1 }\]?

OpenStudy (jcwinner2016):

Ok

OpenStudy (jcwinner2016):

I Think The Answer Is 16/6

OpenStudy (3mar):

then\[\frac{ 3 }{ 4 }-\frac{ 2 }{ 1 }=\] unify the denominators.

OpenStudy (jcwinner2016):

-2 Is Exponent

OpenStudy (3mar):

\[\frac{ 3 }{ 4 }-\frac{ 8 }{ 4 }\] \[= \frac{ 3-8 }{ 4 }=\frac{ -5 }{ 4 }\] This is the final result and no more simplification.

OpenStudy (jcwinner2016):

In My Question Its Says Evaluate

OpenStudy (jcwinner2016):

But Thanks

OpenStudy (3mar):

you just said that now "In My Question Its Says Evaluate". The question does not include "Evaluate"!

OpenStudy (jcwinner2016):

Oh